# David Martínez-Crespo1,2, Cesare Tronci3 1Departamento de Física Teórica, Universidad de Zaragoza, Zaragoza, Spain 2Centro de Astropartículas y Física de Altas Energías, Universidad de Zaragoza, Zaragoza, Spain

### v1

@@ 4 Mixed quantum-classical pure-dephasing systems
where we recall that $0=\eta^{*}({\cal U}^{\dagger}{\rm d}\widehat{\sigma}_{z}{\cal U})={\rm d}^{\widehat{\gamma}}\widehat{\Sigma}_{z}$.
@@
so that the overall expectation values $\int\!f_{0}{\cal Q}\,{\rm d}^{2}z_{0}$ and $\int\!f_{0}{\cal P}\,{\rm d}^{2}z_{0}$ undergo simple harmonic motion. We notice that this again differs from the underlying Heisenberg dynamics above, which instead carries extra terms associated to the backreaction contributions. In the case of the Ehrenfest model, these contributions are absent and one is left with simple harmonic motion also at the level of the underlying Heisenberg dynamics.
@@
where the notation is such that $\{\widehat{\Gamma},{\cal Q}\}^{\gamma}={\cal X}_{Q}\,\lrcorner\,{\rm d}^{\widehat{\gamma}}\widehat{\Gamma}$. We observe that the Koopman model introduces terms orthogonal to $\widehat{\Sigma}_{z}$ in the second equation. These terms are absent in the Ehrenfest model. Moreover, in the latter case, setting $\langle\widehat{\Sigma}_{z}\rangle=0$ yields $\partial_{t}\Omega=0$ while and, as we saw above, the dynamics of the classical variables completely decouples from the quantum dynamics. The situation is very different for the Koopman model, for which the quantum-classical coupling persists at all times due to the backreaction terms.
Evidently, the quantum backreaction produces extra terms in both the velocity and the force which are instead completely overlooked by the Ehrenfest model. Already appearing in the Schrödinger picture [21], this crucial difference makes the Heisenberg dynamics for the Koopman model more realistic and compatible with the fully quantum description. Indeed, in the latter case, the dynamics of the orbital degrees of freedom cannot certainly decouple from the spin motion, thereby leading to a situation very different from the result of the Ehrenfest model and closer to Koopman quantum-classical dynamics.
@@ 5 Conclusions and perspectives
Of the two descriptions considered here, the Ehrenfest model is certainly simpler. While this model is known to fail in reproducing important correlation effects, its underlying geometric structure is particularly rich and has set up the ground for the study of the Koopman model. In the latter, the backreaction term comprising statistical correlations appears explicitly in the total energy, following an analogy with spin-orbit coupling in semi-relativistic mechanics. As already apparent in the Eulerian Schrödinger picture, this backreaction term lifts the difficulty of the equations to a formidable level. Nevertheless, we were able to prove that, for purely (classical) quantum observables, a conservation law for the overall expectation value becomes apparent whenever the observable (Poisson) commutes with the Heisenberg Hamiltonian. This applies to both the models considered here. As for the expectation value of the Hamiltonian itself, we emphasize that this is conserved for the Ehrenfest dynamics, but not for the Koopman case. Evidently, in the latter case this is due to the presence of the backreaction term. This term was shown to be particularly relevant in the simple example of pure dephasing dynamics, in which case the Ehrenfest dynamics becomes trivial for the classical canonical observables and for the vertical component of the quantum spin.
@@ Acknowledgments.
We are grateful to Werner Bauer, Paul Bergold, François Gay-Balmaz, Jesús Clemente-Gallardo, and Delyan Zhelyazov for their comments and remarks on this and related topics. This work was made possible through the support of Grant 62210 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. Financial support by the Leverhulme Research Project Grant RPG-2023-078 is also greatly acknowledged. DMC also acknowledges financial support by Gobierno de Aragón through the grant defined in ORDEN CUS/581/2020.

### v2

@@ 4 Mixed quantum-classical pure-dephasing systems
where we recall that $0=\eta^{*}({\cal U}^{\dagger}{\rm d}\widehat{\sigma}_{z}{\cal U})={\rm d}^{\widehat{\gamma}}\widehat{\Sigma}_{z}$. In this section, we will mostly focus on the Koopman formulation and restrict to consider the Ehrenfest case depending on convenience.
@@
so that the overall expectation values $\int\!f_{0}{\cal Q}\,{\rm d}^{2}z_{0}$ and $\int\!f_{0}{\cal P}\,{\rm d}^{2}z_{0}$ undergo simple harmonic motion. This follows from the fact that $\{{\cal Q},f_{0}\langle i\{{\cal Q},[\hat{\rho}_{0},\widehat{\Sigma}_{z}]\}^{\widehat{\gamma}}\rangle\}^{\widehat{\gamma}}=\Pi({\rm d}{\cal Q},{\rm d}\langle if_{0}\{{\cal Q},[\hat{\rho}_{0},\widehat{\Sigma}_{z}]\}^{\widehat{\gamma}}\rangle)=-\operatorname{div}(\langle if_{0}\{{\cal Q},[\hat{\rho}_{0},\widehat{\Sigma}_{z}]\}^{\widehat{\gamma}}\rangle{\boldsymbol{\cal X}}_{\cal Q})$, and analogously for the corresponding term in the momentum equation. We notice that this expectation dynamics again differs from the underlying Heisenberg evolution of ${\cal Q}$ and ${\cal P}$, which instead carries extra terms associated to the backreaction contributions. In the case of the Ehrenfest model, these contributions are absent and one is left with simple harmonic motion also at the level of the underlying Heisenberg dynamics.
@@
where the notation is such that $\{\widehat{\Gamma},{\cal Q}\}^{\gamma}={\cal X}_{Q}\,\lrcorner\,{\rm d}^{\widehat{\gamma}}\widehat{\Gamma}$. We observe that the Koopman model introduces terms orthogonal to $\widehat{\Sigma}_{z}$ in the second equation. These terms are absent in the Ehrenfest model, for which $\partial_{t}\langle\widehat{\Sigma}_{z}\rangle=0$. Moreover, in the latter case, setting $\langle\widehat{\Sigma}_{z}\rangle=0$ yields $\partial_{t}\Omega=0$ so that, as we saw above, the dynamics of the classical variables completely decouples from the quantum dynamics. The situation is very different for the Koopman model, for which $\langle\widehat{\Sigma}_{z}\rangle$ is not conserved and the quantum-classical coupling persists at all times due to the backreaction terms.
Evidently, the quantum backreaction produces extra terms in both the velocity and the force which are completely overlooked by the Ehrenfest model. Already appearing in the Schrödinger picture [24], this crucial difference makes the Heisenberg dynamics for the Koopman model more realistic and compatible with the fully quantum description. Indeed, in the latter case, the dynamics of the orbital degrees of freedom cannot decouple from the spin motion, thereby leading to a situation very different from the result of the Ehrenfest model and closer to Koopman quantum-classical dynamics.
@@ 5 Conclusions and perspectives
Of the two descriptions considered here, the Ehrenfest model is certainly simpler. While this model is known to fail in reproducing important correlation effects, its underlying geometric structure is particularly rich and has set up the ground for the study of the Koopman model. In the latter, the backreaction term comprising statistical correlations appears explicitly in the total energy, following an analogy with spin-orbit coupling in semi-relativistic mechanics. As already apparent in the Eulerian Schrödinger picture, this backreaction term lifts the difficulty of the equations to a formidable level. Nevertheless, we were able to prove that, for purely (classical) quantum observables, a conservation law for the overall expectation value becomes apparent whenever the observable (Poisson) commutes with the Heisenberg Hamiltonian. This applies to both the models considered here. As for the local expectation $\langle\widehat{\cal H}\rangle$ of the Hamiltonian itself, we emphasize that this is conserved for the Ehrenfest dynamics, but not for the Koopman case. Evidently, in the latter case this is due to the presence of the backreaction term. This term was shown to be particularly relevant in the simple example of pure dephasing dynamics, in which case the Ehrenfest dynamics becomes trivial for the classical canonical observables and for the vertical component of the quantum spin. We remind the reader that the local expectation $\langle\widehat{\cal H}\rangle$ must not be confused with the total energy ${\int}f_{0}\langle\widehat{\cal H}\rangle{\rm d}^{2}z_{0}$ which is consistently conserved in both the Ehrenfest model. Likewise, the Koopman model conserves the energy functional ${\int}f_{0}\langle\widehat{\cal H}+i\hbar[\nabla^{\widehat{\gamma}}\hat{\rho}_{0},{\boldsymbol{\mathcal{X}}}_{{\widehat{\mathcal{H}}}}]/2\rangle{\rm d}^{2}z_{0}$, whose second term generates quantum-classical correlations.
@@
We conclude by emphasizing that, despite the level of complexity in the Heisenberg dynamics, both the Ehrenfest and Koopman models possess convenient particle closures [49] that are formulated combining their underlying variational principle with the momentum map structures associated to their geometry. In the Koopman case, the numerical implementation has been successfully benchmarked in [7]. This is an example in which the complexity of a PDE system may be conveniently overcome by resorting to its underlying geometric structure. In this sense, while the hybrid Schrödinger picture remains considerably simpler, the geometric structure of the corresponding Heisenberg dynamics presented here may lead to new directions to be unfolded in the future. One possible example is to transfer the Heisenberg picture –or any of its extensions proposed above– to the level of the particle closure in such a way to tackle specific problems where the Schrödinger picture is less convenient.
Data Availability Statement:
No Data associated in the manuscript
@@ Acknowledgments.
We are grateful to Werner Bauer, Paul Bergold, Jesús Clemente-Gallardo, François Gay-Balmaz, Darryl Holm, Alberto Ibort, and Delyan Zhelyazov for their comments and remarks on this and related topics. This work was made possible through the support of Grant 62210 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. Financial support by the Leverhulme Research Project Grant RPG-2023-078 is also greatly acknowledged. DMC also acknowledges financial support by Gobierno de Aragón through the grant defined in ORDEN CUS/581/2020 as well as the support of Grant PID2021-123251NB-I00 funded by MCIN/AEI/10.13039/ 501100011033 and by the European Union, and of Grant E48-23R funded by Government of Aragon..